OpenStudy (anonymous):

How do I get started on this problem? 2x/x^2-4 = 4/x^2-4 - 3/x+2

OpenStudy (aum):

Is this the problem? \[ \frac{2x}{x^2-4} = \frac{4}{x^2-4} - \frac{3}{x+2} \]

OpenStudy (aum):

If so, multiply throughout by \((x^2-4)\) and simplify. Then solve for x.